OpenStudy (anonymous):

Solve. 4 + x/7=2 A) -14 B) 10 C) 12 D) 42

OpenStudy (anonymous):

nvm i got it ty

OpenStudy (anonymous):

4 + x/7 = 2 4 - 4 + x/7 = 2 - 4 x/7 = -2 x/7 x 7 = -2 x 7 x = -14

OpenStudy (anonymous):

-14

OpenStudy (anonymous):

:) Thanks!